# Bipartite and tripartite entanglement in pure dephasing relativistic spin-boson model

### v1

@@ II.2 UDW model as a time-independent RSB model
    It allows for analysis of relativistic causality in a manner closer to the standard quantum optics. For example, it is much simpler to specify how long the interaction has to occur before the detectors can effectively signal to one another by checking that $t\gtrsim L$, where $L$ is the separation of their COM. This is particularly useful for one of our results, where we will show that significant entanglement is only generated “deep in the light cone interior” (c.f. Section III).
For completeness, we will extend the regularity requirements for single detectors in [29] to $N$ detector settings. These will be relevant for a better understanding of how the entanglement dynamics of the two initially separable detectors depend on the mass of the field and the spacetime dimensions. This goes by the name strong Huygens principle, which we briefly discuss below.
@@ II.3 Strong Huygens principle
One of our goals is to understand the impact of the mass and spacetime dimensions on the entanglement dynamics between two detectors in the non-perturbative regime, since these modify the signaling contribution to the dynamics. The perturbative analysis has been covered in the literature [41, 64, 65].
Roughly speaking, a field $\phi$ is said to satisfy the strong Huygens principle when its Green’s function has distributional support only on the light cone [38]. Intuitively, this means that an impulsive light emanated from an emitter can be captured only by a lightlike separated receiver. When support extends inside the light cone, we say the strong Huygens principle is violated, and a signal can be captured by a timelike receiver. By construction, massive fields violate the strong Huygens principle as massive excitations travel at subluminal speed. The simplest example of the strong Huygens principle is the massless scalar field in $(3+1)$-dimensional Minkowski spacetime, where in the inertial coordinates the causal propagator is given by
@@
Clearly, the field obeys the strong Huygens principle by virtue of the fact that the Dirac delta function has no support except when $\Delta t=\pm|\Delta{\bm{x}}|$, i.e., two points are null-separated. It is also known that in $(n+1)$-dimensional flat spacetime, the wave equation $\square\phi=0$ fails the strong Huygens principle for odd spacetime dimension ($n$ even) and when $n=1,2$. In even spacetime dimensions the strong Huygens principle is satisfied (see [41, 58] for some explicit computations).
Remarkably, some generic results about the violation of the strong Huygens principle can be obtained for curved spacetimes [40, 39]. Some intuition can be gleaned from the Klein-Gordon equation in (1): if we consider a generic non-minimal coupling parameter $\xi>0$, then massless fields coupled to gravity acquire effective mass due to the Ricci scalar contribution. Indeed, for constant-curvature geometry, this effectively gives a mass parameter $m_{\text{eff}}\coloneqq\sqrt{\xi R}$, so the Klein-Gordon equation is equivalent to that of a massive field minimally coupled to gravity ($\xi=0$). The importance of the strong Huygens principle is that it affects the propagation of information: indeed, massless fields that violate the strong Huygens principle can signal to the interior of the light cone as the Green’s function between two timelike-separated events is non-vanishing. The violation of the strong Huygens principle has been shown, using perturbative techniques, to impact how two UDW detectors can be entangled when they are in causal contact [41, 64] and how they can be exploited to send timelike signals using massless fields [66, 67].
@@ II.4 Regularity conditions
Before we proceed with analysis of the bipartite entanglement dynamics, we first take a detour to establish some regularity requirements for the RSB model for $N$-emitters. This will justify in part the use of our model instead of the usual interaction picture, and also why it will be useful to work with covariant IR-cutoff $m\ll 1$ (in an appropriate unit) for dealing with massless case. Readers interested only in the physical results pertaining entanglement dynamics of the two detectors can skip to Subsection III. In order to keep the discussion fairly accessible, we will try to frame our results in rigorous terms but with fairly physicist-angled explanation.
@@
By ‘zero mode’ we mean the field mode that is not associated with the 1-particle sector of the Fock space, i.e., it behaves like a free particle in quantum mechanics [68, 69, 46, 70] and hence does not admit a Fock representation. This complicates the choice of the vacuum state on which the full Hilbert space can be built (see, e.g., [68, 69] for simple examples on Einstein cylinder), so we will avoid them for simplicity as they can be dealt only case-by-case.
@@
Although it is not useful for studying weakly-coupling phenomenon such as the Unruh effect, the pure-dephasing regime in Assumption 1 is important as they can still exhibit rich dynamics and they enable rigorous results. More technically, the pure-dephasing model can be viewed as a $C^{*}$-dynamical system, which roughly speaking means that the dynamics can be implemented by an automorphism $\alpha^{0}_{t}:\mathcal{A}\to\mathcal{A}$ where $\mathcal{A}$ is the joint algebra of observables of the detector-field system, independently of the Hilbert space representations. This is to be contrasted with the nonzero gap case, where the corresponding automorphism can only be rigorously defined as a pseudodynamics [48], i.e., one has to work with an implementation of $\alpha_{t}^{\Omega}:\mathcal{B}(\mathcal{H}_{\omega})\to\mathcal{B}(\mathcal{H}_{\omega})$ in a particular Hilbert space representation. Morally speaking, this is due to the fact that for $\Omega\neq 0$, there is no guarantee that $\alpha_{t}^{\Omega}$ will not “bring us outside the Hilbert space” and give rise to divergences. The best one can do is to do a “small $\Omega$” perturbative calculation about the pure-dephasing as the perturbation is bounded [31]. Treating the interaction Hamiltonian as a perturbation to the free Hamiltonian is much harder due to unbounded nature of the field operator (see, e.g., [49, 50]).
@@
Apart from the issue of mathematical rigor, the main physical reason for choosing Assumption 1 is for the sake of having rigorous results even for the strong coupling regime where the coupling constants $\lambda_{j}$ for each detector can be chosen to be arbitrarily large. In the original UDW case, it was shown that if the interaction Hamiltonian of the UDW model (15) has spacelike supports, then the reduced density matrix of the two detectors after interaction is separable [22]. The key, of course, is spacelike separation as we know that generic interacting quantum systems that are in causal contact will generate some entanglement, and the same gapless model was shown to generate entanglement when they are in causal contact [37].
@@
As we will show in Section III, one of our contribution will be to show that not only does the RSB model satisfy the no-go theorem, but that the detectors can only get entangled very deep into the light cone. In other words, even though we can in principle make the final state arbitrarily close to the maximally entangled state, it takes much longer than the light-crossing time between them. To set up the stage for these results, we need more assumptions along the lines used for the single detector case in [29]. These additional assumptions require us to first study the ground state properties of the full detector-field system.
The first result is to guarantee that the ground state energy is bounded from below, assuming the existence of a ground state from the GNS construction. [^7] We are following the construction in [31]: there, one assumes first the existence of a ground state in the form of some vector |$\ket{\psi}\in\mathcal{H}_{\omega}\cong\mathbb{C}^{2}\otimes\mathfrak{F}(\mathcal{H})$, and then find the ground state energy. The inadequacy of this construction comes from the issue of whether $\mathfrak{F}(\mathcal{H})$ is the non-interacting Fock space, in which case one has to construct the state by, e.g., taking zero-temperature limit of some thermal state. From the calculational point of view, it will turn out that the ground state energy will be the same either way as we can just view |$\ket{\psi}$ as a GNS vector.
@@
The proof is given in Appendix A for completeness.
Proposition 2.
The proof is a simple adaptation of [31] except that we need to consider all choices of signs $\bm{s}$ and that the integral $R_{2}$ is with respect to the “combined smearing function” $\bm{s}\cdot\bm{F_{k}}$. In particular, we see that the model exhibits IR divergences (production of an infinite number of soft bosons in the thermodynamic limit) if at least one detector is not coupled to the field with a sufficiently IR-regular coupling function $F^{(j)}_{\bm{k}}$.
As argued in [29], this suggests that modelling a physical system with the RSB or UDW framework requires careful control of both UV and IR behavior. The UV regularity is automatically taken care of by giving each detector a finite size (effectively serving as a UV cutoff), but the IR case is more subtle. If the field is massive or gains effective mass through, say, coupling to the (classical) gravitational field, then the model is also IR-regular. In contrast, when there are soft bosons, such as (3+1)-dimensional Minkowski spacetime with massless field, the key insight to be learnt from [31, 32, 30, 49, 50] is to consider a different Hilbert space representation that is not unitarily equivalent to the vacuum representation: indeed, the Araki-Woods representation and the thermal KMS representation suffice for the situation at hand [71, 72] (see [62] for very recent analysis on KMS states). At the level of correlation functions, the vacuum correlators can be computed by taking the zero-temperature limit of the thermal correlators.
Since in this work we are going to focus on entanglement dynamics, the situation involving the infinitely many soft boson productions will not be of concern to us. Indeed, perturbatively one can check (following the strategy in [41]) that very light scalar field ($m\approx 0$) leads to bipartite entanglement between the detectors that is indistinguishable compared to the massless case. Thus we are going to assume this:
Assumption 2.
In the context of our results, we will consider the field to be massless for parameter $m\sigma\ll 10^{-11}$ where $\sigma$ is the effective size of the detector. Crucially, this implies that the RSB model is both UV- and IR-regular in the sense above, it follows that the non-interacting Hilbert space is unitarily equivalent to the joint interacting Hilbert space, i.e., the interacting Hilbert space is $\mathbb{C}^{2}\otimes\mathbb{C}^{2}\otimes\mathfrak{F}(\mathcal{H})$ where $\mathfrak{F}(\mathcal{H})$ is the Fock space of the free field theory. In the subsequent analyses we will make these assumptions for simplicity, noting that one can always pass to the correct Hilbert space representation when the IR regularity needs to be dropped [29, 32, 31].
III Entanglement dynamics between two detectors
[…]
  1. (i)
    Perturbatively, spacelike-separated detectors with interaction Hamiltonian not commuting with their free Hamiltonian can be entangled starting from initially separable state. This is the well-known entanglement harvesting protocol [73, 74, 75, 76, 77], and has been studied extensively in various scenarios (curved spacetimes, different detector models, etc.).
  2. (ii)
    Perturbatively, detectors can be entangled also when the interaction is causally connected [73], however this entanglement can be independent of the vacuum entanglement of the field state depending on the wave property known as the strong Huygens principle [41, 65, 78].
  3. (iii)
    Non-perturbatively, it is known that delta-coupled detectors cannot get entangled at spacelike separation if it is “simple-generated” [22] (i.e., each detector interacts instantaneously once [27]), but can get entangled otherwise.
  4. (iv)
    Non-perturbatively, gapless detector models cannot extract entanglement at spacelike separation, based on arguments about entanglement-breaking channels [22, 79]. Consequently, initially separable detector remains separable after interaction.
  5. (v)
    Non-perturbatively, it was shown in [37] for massless fields in (3+1)D Minkowski spacetime that gapless models can get entangled if interaction time is long enough for detectors to be in causal contact. [^8] It was also shown in [37] that if the coupling is also sufficiently weak, the two-detector dynamics can be approximated in some sense by an entangling unitary. The latter is essentially the same spirit as the well-known semiclassical optics calculation where the electromagnetic field is not quantized [80, 81]. This analysis was motivated partially to better understand gravity-induced entanglement setups [82].
There are countless results in the literature involving similar setups but different detector trajectories, boundary conditions, spacetime geometries, initial states, etc. that are only tangentially relevant for our purposes, so we do not aim for an exhaustive list.
Our initial goal is to complement the analysis in the literature concerning (ii), (iv) and (v). That is, we carry out a fully non-perturbative analysis of entanglement dynamics between two gapless detectors that are in causal contact in scenarios where the strong Huygens principle is violated, i.e., accounting for the effects of mass and spacetime dimensions. As we show below, pursuing this goal has a good payoff: it turns out that the entanglement dynamics is actually richer and non-trivial once we include mass and spacetime dimension as tunable parameter.
We first recall that the original interest in the so-called entanglement harvesting protocol is fundamental in nature, since it was to see whether vacuum entanglement can be detected at all [73, 74, 77, 75, 76]. For this reason, it is immaterial that the amount of entanglement is extremely small or whether the calculations are only perturbative in nature, as the main purpose is as a sort of “witness” for vacuum entanglement in the field. From a theoretical standpoint, there is nothing very surprising about the protocol itself: what is non-trivial is whether it is detectable in practice. Several experimental proposals have been given, from the early proposal involving linear ion trap [83] to the recent superconducting circuits [84], with recent experimental realization via electro-optics sampling experiment [85].
On the other hand, if the goal is to generate entanglement, the quality matters and it is somewhat immaterial that the entanglement is extracted from the vacuum of the field (which would be very small in general). Indeed, physical intuition tells us that it is generically unsurprising for two quantum systems A and B to get entangled in general if we allow them to have direct interactions and one has to engineer the interactions and the initial states to avoid getting entangled. The same is true if we allow induced interactions, i.e., the two systems interact with a common third party C that induces effective interactions between the two. Essentially, C mediates the interaction by propagating signals from A to B (it could be phonons, photons, etc.). Consequently, the physically relevant question is how the entanglement is generated and how good it can be as a function of parameters. This necessitates non-perturbative approach, as perturbation theory only allows perturbatively small amount of entanglement.
[…]
In Fig. 1 we plot the negativity of entanglement $\mathcal{N}$ as a function of separation $L$ between the detectors’ center of mass and the duration of interaction $t$ for the weak coupling regime $\widetilde{\lambda}=0.01$, with a) showing the massless regime and b) showing the massive regime. In both massless and massive cases, we see that the UDW-type coupling can be used to generate bipartite states close to being maximally entangled. [^9] For the massless case, the fact that we can generate significant amount of entanglement can already be inferred from [37] with suitable adjustment and rescaling (to account, for instance, the Gaussian switching). However, perhaps surprisingly, both Fig. 1a-i) and b-i) show that the light cone (white dashed line) has a much gentler slope than the black boundary which marks the beginning of nonzero entanglement — for convenience we call these “entanglement cone”. In other words, significant amount of entanglement can only be generated very deep into the light cone. This is to be contrasted with the standard perturbative result for the UDW model, where entanglement peaks in general at the light cone [41]. The entanglement cone is not linear in the $t-x$ plane, and we see that increasing the mass makes the entanglement cone steeper (i.e., it takes much longer time to reach maximum entanglement).
In Fig. 2 we plot the same scenario in (3+1)D for the strong coupling regime $\widetilde{\lambda}=1$. In the strong coupling regime, we observe that the entanglement cone is unchanged compared to the weak coupling, however the maximum amount of entanglement is much lower in both massless and massive cases. In fact, we observe that not only does the massless entanglement decay in time, but its maximum is lower than that for the massive field. In other words, mass enhances the amount of entanglement generated for the two detectors. Note that because the RSB model used here is gapless, this cannot be explained by any resonance-type argument (which works in the perturbative regime as in [65]). Importantly, this is not contradicting any intuition about “exponential decay” associated with mass: mass contributes to the exponential decay of correlations as a function of separation $L$, but not as a function of temporal duration $t$ — this can be checked, for instance, by looking at the asymptotic behaviour of the Wightman two-point functions in the large-$t$ or large-$L$ limit.
At this stage, it is unclear if this phenomenon is tied to the strong Huygens principle. If it were, then the same result should be expected by varying the spacetime dimensions rather than mass. We show that this is intrinsically due to mass rather than the violation of the strong Huygens principle in Figure 3 for the weak coupling regime. Here we can see clearer in the rightmost figures a-ii)-c-ii) that mass generically improves entanglement generation, in that it always produces at least as much entanglement as the massless case, as is the case in (4+1)D and (5+1)D. In (2+1)D, the separation between massless and massive case is huge: we see that the massless field can only generate half as large negativity as the massive case. What is true in general, however, is that massive fields attain maximum entanglement much slower than the massless field. We note in passing that the entanglement cone is steeper as the spacetime dimension is increased — this is itself not surprising, as this can be explained by the fact that the Wightman two-point functions has stronger decay (in both space and time directions) in higher dimensions. It implies that generating entanglement for a given separation is harder in higher spacetime dimensions.
Now that we singled out mass as the factor that contributes to the enhancement of entanglement between the two detectors, it is useful to see how entanglement pattern changes with mass as we fix either spatial separation or duration of interaction, as shown in Figure 4. Here we see that the properties of spatial and temporal correlations of the field are indeed asymmetric and influences the negativity differently. [^10] Note that the asymmetry in the way correlations decay in the spatial and temporal directions is not unique to relativistic scalar fields: it is a well-known phenomenon in quantum many-body and condensed matter physics, which underlies why quantum many-body dynamics is a hard problem even for systems with exponentially decaying spatial correlations. Notice that in Figures (a-i) and (b-i), we understand in what sense massless fields are more helpful: the larger the mass, the faster the negativity vanishes as a function of $L$, hence massless fields has “larger reach” in spatial separation as negativity vanishes at larger separation. Figures (a-ii) and (b-ii) complement our observations earlier: mass generically improves the maximum amount of entanglement generated on the detectors, in this case for the strong coupling regime relative to the weak-coupling one.
We summarize the main results as follows:
  1. (i)
    Non-perturbatively, gapless models can indeed generate entanglement between the two detectors, but they generically do so very deep into the light cone. This is in contrast to the standard perturbative result where entanglement is generated starting at the light cone (in other words, light cone and entanglement cone coincides) [41]. This is a stronger result than the no-go theorem in [22], which only proves that entanglement cannot be generated at spacelike-separation, since we showed that it is not enough that their interactions are in causal contact.
  2. (ii)
    Mass can improve entanglement generated relative to massless case in terms of the amount of negativity, sometimes approaching maximal entanglement $\mathcal{N}\approx 0.5$ when massless field cannot. This is independent of the violation of the strong Huygens property. However, in general, massive fields only achieve maximum negativity at much longer interaction time, consistent with the fact that the entanglement cone is steeper with increasing mass.
  3. (iii)
    There is clear asymmetry between decay of correlations in spatial and temporal directions, which manifests as the difference in how negativity of the two detectors decay as a function of time and spatial separation for different mass parameters.
IV Entanglement dynamics between three detectors
In this section, we study the entanglement dynamics between three gapless RSB-type UDW detectors interacting with a scalar field initialized in the uncorrelated ground state of the non-interacting theory ($\lambda_{j}=0$) given by
[…]
For simplicity, the three detectors are assumed to be static and placed symmetrically at the vertices of an equilateral triangle, each separated by a distance $L$ from one another [42]. This simplification is also helpful when trying to isolate the physics associated with the detector-field dynamics from the non-uniformity introduced by the asymmetric placement of the detectors.
Due to the relatively painful amount of computations, tripartite dynamics in the standard UDW model is already sparse. The relevant known results so far are:
  1. (i)
    Perturbatively, it was shown in Minkowski spacetime that three UDW detectors can extract tripartite entanglement at spacelike separation in various spatial configurations, not only the equilateral one [42]. This is also demonstrated for (2+1)D BTZ black hole spacetime [43]. The authors used $\pi$-tangle [86] that serves as a lower bound for the amount of genuine tripartite entanglement in mixed states of three qubits. It is noteworthy that $\pi$-tangle is not an entanglement measure on its own. There were some earlier studies that show W-type entanglement in entanglement harvesting-type setups (i.e., for spacelike-separated detectors) [87].
  2. (ii)
    Non-perturbatively, it was shown for Minkowski [28] that three UDW detectors can get entangled for delta-coupled detector model, also using $\pi$-tangle as a ‘witness’ of genuine tripartite entanglement. They show that there exist parameter regimes for which the tripartite entanglement is GHZ-type.
As before, we do not aim for an exhaustive list. From these, we believe that the long-time regime involving the gapless model has not been fully understood prior to this work. Let us remark that the regularity requirements for the three-detector RSB model follows directly from the general case in Section II.4 and Appendix A.
It is well-known that constructing entanglement measures for multipartite entanglement is highly non-trivial. Even for the case of three qubits, it is known that it is not possible to have a single entanglement measure that unambiguously tells whether one state is more entangled than another [88]. This is related to the fact that for multipartite entanglement one can have inequivalent but genuine multipartite entanglement (GME): the simplest examples being the GHZ and W states, which are in some sense “maximally entangled” (see, e.g., [89]). For our purposes, this issue is complicated by the fact that if we consider the RSB model with three detectors, then the reduced density matrix of the detector is mixed: for mixed states, constructing faithful entanglement measure is generically difficult, and none of the convex-roof extensions (if at all) are easy to compute for mixed states.
In order to analyse the tripartite entanglement, we follow [42, 28, 43] and consider the $\pi$-tangle, defined as
Here,
$\mathcal{N}_{\mathrm{A(BC)}}$ $\coloneqq\frac{||\rho_{\mathrm{ABC}}^{\intercal_{\mathrm{A}}}||_{1}-1}{2}\,,$ (27a)
$\mathcal{N}_{\mathrm{A(B)}}$ $\coloneqq\frac{||\rho_{\mathrm{AB}}^{\intercal_{\mathrm{A}}}||_{1}-1}{2}\,,$ (27b)
$\mathcal{N}_{\mathrm{A(C)}}$ $\coloneqq\frac{||\rho_{\mathrm{AC}}^{\intercal_{\mathrm{A}}}||_{1}-1}{2}\,,$ (27c)
are the negativities of the three-detector system. The $\pi$-tangle quantifies the tripartite entanglement if the joint state of the detectors $\rho_{\mathrm{ABC}}$ is pure due to the Coffman-Kundu-Wootters inequality for negativity, $\pi_{i}\geq 0$, $\forall i\in\{\text{A},\text{B},\text{C}\}$ [90, 86]. Unfortunately, this property fails when $\rho_{\mathrm{ABC}}$ is mixed, and the $\pi$-tangle serves only as a lower bound for the tripartite entanglement. This suggests that the $\pi$-tangle for mixed states could be zero or negative, even if there exists tripartite entanglement among the detectors. Nevertheless, tripartite entanglement is guaranteed at least when the $\pi$-tangle is positive. For this reason, in what follows we take $\max\{\pi,0\}$ and examine the parameter space with positive $\pi$-tangle.
  1. (i)
    There exists a regime in the parameter space for which the three detectors have genuine tripartite entanglement of the “GHZ type”, in the sense that for any bipartition has zero entanglement. However, this regime is generically very narrow and has very small $\pi$-tangle, thus it is difficult to certify that the tripartite entanglement is significant.
  2. (ii)
    A large portion of the parameter space allows for tripartite entanglement such that any bipartition has nonzero entanglement. Similar to the bipartite case, mass of the field tends to amplify the $\pi$-tangle, but there is no natural way to discern the nature of the tripartite entanglement.
We will provide arguments for why this is the case and put the existing known results in perspectives.
Our argument is based on a numerical analysis in Figure 5 for both weak and strong coupling as well as massless and massive fields. The white dashed line represents the light cone. What we show here is that on the one hand, the black region below the red line in Figure 5 corresponds to the case for which the three detectors have genuine tripartite entanglement of the “GHZ type”, in the sense that for any bipartition has zero entanglement. However, this region is generically very narrow, thus one has to fine-tune the arrangement of the detectors to obtain this. Furthermore, since the $\pi$-tangle is very small, it is difficult to certify that this genuine tripartite entanglement is actually significant. On the other hand, the area bounded by the green dots above the red line shows the regime that allows for tripartite entanglement with nonzero bipartite entanglement. We also observe that mass of the field tends to amplify the $\pi$-tangle, however there is no natural way to obtain precise information about the nature of the tripartite entanglement. Indeed, we checked that none of the usual tripartite entanglement witness for mixed states works [91], as we verified them numerically and all of them are inconclusive (i.e., they say nothing about the tripartite entanglement content). These results seem to suggest that there are significant limitations in using gapless UDW or RSB-type models to learn about both intrinsic multipartite entanglement and multipartite entanglement generation in QFT.
A natural question to ask is how this result is compatible with the observations in [28], which essentially claim that there is genuine tripartite entanglement of GHZ type in Minkowski spacetime and BTZ spacetime for the vacuum state of a massless scalar field with delta-coupled model. Strictly speaking, while there are some similarities between delta-coupled model and the gapless model [27], the results in [28] is not ideal as they showed genuine tripartite entanglement only for very small detector separation, effectively with significant overlap of their Gaussian smearing functions. In this regime, one can question the validity of the standard UDW model. This is because very closely-spaced atoms may experience interatomic interactions that the UDW model we considered here does not capture. In this sense, our results are more reliable as they do not rely on significant overlap of the spatial smearing and still displays genuine tripartite entanglement of GHZ type.
V Discussion and outlook
In this work we have studied non-perturbatively the entanglement generation between two emitters in an exactly solvable relativistic variant of the spin-boson model, equivalent to the time-independent formulation of the (gapless) Unruh-DeWitt model. This formulation allows for rigorous basis of the regularity of the model in the UV and in the IR and allows slightly cleaner analysis of the causal behaviour of the interactions.
We showed that (i) (highly) entangled states of the two emitters (detectors) require interactions very deep into the light cone, strengthening the no-go result of [22] that says there can be no entanglement for spacelike-separated interactions; (ii) the mass of the field generically improves the maximum amount of entanglement generated on the detectors at the expense of longer interaction times and furthermore, we argued that this is not fundamentally due to the strong Huygens principle; (iii) unlike the bipartite case where it is possible to generate close to maximally entangled states via the spin-boson interactions, the generation of genuine tripartite entanglement for sufficiently separated detectors is non-perturbatively hard even at sufficiently long times, refining the delta-coupling result in [28]. Result (iii), in particular, suggests that probing the multipartite entanglement of a relativistic quantum field non-perturbatively requires either different probe-based techniques or variants of the UDW model.
In a way, our results suggest that despite the non-perturbative and simple nature of the model, the gapless variant of the UDW model is not good enough once we are interested in obtaining more precise information about multipartite entanglement, whether to study the intrinsic multipartite entanglement in the field or the ability for relativistic quantum fields to mediate entanglement generation. From the quantum-optical standpoint, it means that this model is not very useful for generating highly-entangled multipartite entangled states even if we allow causal contact, and since this model is unable to extract entanglement from the vacuum state at spacelike separation, we can think of the gapless model as being suitable, for instance, in studying classical communication settings (see, e.g., [26]). That said, our work closes the (remaining) gap for the two-detector scenario: we showed that causal contact enables entanglement generation in a maximal sense, but this requires the interaction time to be much larger than the spatial separation of the detectors, in contrast to the known perturbative results [41].
Three natural questions arise from our work. First, is it possible to consider a variant of the setup or the model in such a way that we can probe non-perturbatively multipartite entanglement of the vacuum state of the field in a useful way? Presumably, there are operator-algebraic approach to studying multipartite entanglement [92] of the QFT directly, and this should ideally be complementary. Second, can multiple delta-coupled detectors be used more effectively to study entanglement in QFT? The first use case was in [22], and recently multi-delta coupling has been used in single-detector settings [93, 24]. These quickly become unwieldy for large number of couplings and detectors, but in principle computable. Last but not least, one way to resolve the ambiguity regarding the nature and “strength” of the tripartite entanglement is by operational means, i.e., by identifying protocols (e.g., teleportation in or steering in the bipartite case) that requires certain kinds and amount of tripartite entanglement to succeed and see if the tripartite state produced via the RSB interaction works for these protocols. We leave these questions for the future.
Acknowledgment
Appendix A Regularity for N-partite RSB model
Proof of Proposition 1.
@@ Appendix A Regularity for N-partite RSB model
where $\bm{\Delta}\equiv[\Delta_{\mathrm{A}},\Delta_{\mathrm{B}},\cdots]^{\intercal}$ is a vector of $\Delta_{j}$, and $\bm{F_{k}}\equiv[F_{\bm{k}}^{(\text{A})},F_{\bm{k}}^{(\text{B})},\ldots]^{\intercal}$ is a vector of coupling functions. As an example, consider two detectors A and B (i.e., when $N=2$). In this case, we have |$\ket{\bm{s}}\in\{\ket{+_{\mathrm{A}},+_{\mathrm{B}}},\ket{+_{\mathrm{A}},-_{\mathrm{B}}},\ket{-_{\mathrm{A}},+_{\mathrm{B}}},\ket{-_{\mathrm{A}},-_{\mathrm{B}}}\}$, and so the gapless RSB Hamiltonian reads
@@
This suggests that the ground state of $\hat{H}_{\phi,0,\bm{s}}$ is given by the coherent state |$\ket{\alpha_{\bm{s}}}\equiv\hat{D}^{\dagger}(\alpha_{\bm{s}})\ket{0}$, where |$\ket{0}$ is the Fock vacuum of the free Hamiltonian $\hat{H}_{\phi,0}$. Therefore, defining |$\ket{\psi_{0}^{(\bm{s})}}\coloneqq\ket{\bm{s}}\otimes\ket{\alpha_{\bm{s}}}$, $\bm{s}\in\{\pm\}^{N}$, we obtain

### v3

@@ II.2 UDW model as a time-independent RSB model
    It allows for analysis of relativistic causality in a manner closer to the standard quantum optics. For example, it is much simpler to specify how long the interaction has to occur before the detectors can effectively signal to one another by checking that $t\gtrsim L$, where $L$ is the separation of their COM. This is particularly useful for one of our results, where we will show that significant entanglement is only generated “deep in the light cone interior” (c.f., Sec. III).
For completeness, we will extend the regularity requirements for single detectors in [Tjoa2024interacting] to $N$ detector settings. These will be relevant for a better understanding of how the entanglement dynamics of the two initially separable detectors depend on the mass of the field and the spacetime dimensions. This goes by the name strong Huygens principle, which we briefly discuss below.
@@ II.3 Strong Huygens principle
One of our goals is to understand the impact of the mass and spacetime dimensions on the entanglement dynamics between two detectors in the nonperturbative regime, since these modify the signaling contribution to the dynamics. The perturbative analysis has been covered in the literature [tjoa2021harvesting, Casals2020commBH, hector2022optimize].
Roughly speaking, a field $\phi$ is said to satisfy the strong Huygens principle when its Green’s function has distributional support only on the light cone [McLenaghan1974huygens]. Intuitively, this means that an impulsive light emanated from an emitter can be captured only by a lightlike separated receiver. When support extends inside the light cone, we say the strong Huygens principle is violated, and a signal can be captured by a timelike receiver. By construction, massive fields violate the strong Huygens principle as massive excitations travel at subluminal speed. The simplest example of the strong Huygens principle is the massless scalar field in $(3+1)$-dimensional Minkowski spacetime, where in the inertial coordinates the causal propagator is given by
@@
Clearly, the field obeys the strong Huygens principle by virtue of the fact that the Dirac delta function has no support except when $\Delta t=\pm|\Delta{\bm{x}}|$, i.e., two points are null-separated. It is also known that in $(n+1)$-dimensional flat spacetime, the wave equation $\square\phi=0$ fails the strong Huygens principle for odd spacetime dimension ($n$ even) and when $n=1,2$. In even spacetime dimensions, the strong Huygens principle is satisfied (see [tjoa2021harvesting, Causality2015Eduardo] for some explicit computations).
Remarkably, some generic results about the violation of the strong Huygens principle can be obtained for curved spacetimes [Faraoni2019huygens, Sonego1992huygenscurved]. Some intuition can be gleaned from the Klein-Gordon equation in (1): if we consider a generic nonminimal coupling parameter $\xi>0$, then massless fields coupled to gravity acquire effective mass due to the Ricci scalar contribution. Indeed, for constant-curvature geometry, this effectively gives a mass parameter $m_{\text{eff}}\coloneqq\sqrt{\xi R}$, so the Klein-Gordon equation is equivalent to that of a massive field minimally coupled to gravity ($\xi=0$). The importance of the strong Huygens principle is that it affects the propagation of information: indeed, massless fields that violate the strong Huygens principle can signal to the interior of the light cone as the Green’s function between two timelike-separated events is nonvanishing. The violation of the strong Huygens principle has been shown, using perturbative techniques, to impact how two UDW detectors can be entangled when they are in causal contact [tjoa2021harvesting, Casals2020commBH] and how they can be exploited to send timelike signals using massless fields [Blasco2015Huygens, Blasco2016broadcast].
@@ II.4 Regularity conditions
Before we proceed with the analysis of the bipartite entanglement dynamics, we first take a detour to establish some regularity requirements for the RSB model for $N$ emitters. This will justify in part the use of our model instead of the usual interaction picture, and also explain why it will be useful to work with a covariant IR-cutoff $m\ll 1$ (in an appropriate unit) for dealing with the massless case. Readers interested only in the physical results pertaining to the entanglement dynamics of the two detectors can skip to Sec. III. In order to keep the discussion fairly accessible, we will try to frame our results in rigorous terms but with a fairly physicist-angled explanation.
@@
By zero mode we mean the field mode that is not associated with the 1-particle sector of the Fock space, i.e., it behaves like a free particle in quantum mechanics [Tjoa2019zeromode, EMM2014zeromode, KayWald1991theorems, folacci1987desitter] and hence does not admit a Fock representation. This complicates the choice of the vacuum state on which the full Hilbert space can be built (see, e.g., [Tjoa2019zeromode, EMM2014zeromode] for simple examples on Einstein cylinder), so we will avoid them for simplicity as they can be dealt only case-by-case.
@@
Although it is not useful for studying weakly coupling phenomenon such as the Unruh effect, the pure-dephasing regime in Assumption 1 is important as they can still exhibit rich dynamics and they enable rigorous results. More technically, the pure-dephasing model can be viewed as a $C^{*}$-dynamical system, which roughly speaking means that the dynamics can be implemented by an automorphism $\alpha^{0}_{t}:\mathcal{A}\to\mathcal{A}$ where $\mathcal{A}$ is the joint algebra of observables of the detector-field system, independently of the Hilbert space representations. This is to be contrasted with the nonzero gap case, where the corresponding automorphism can only be rigorously defined as a pseudodynamics [bratteli2002operatorv2], i.e., one has to work with an implementation of $\alpha_{t}^{\Omega}:\mathcal{B}(\mathcal{H}_{\omega})\to\mathcal{B}(\mathcal{H}_{\omega})$ in a particular Hilbert space representation. Morally speaking, this is due to the fact that for $\Omega\neq 0$, there is no guarantee that $\alpha_{t}^{\Omega}$ will not “bring us outside the Hilbert space” and give rise to divergences. The best one can do is to do a “small $\Omega$” perturbative calculation about the pure-dephasing as the perturbation is bounded [spohn1989spinboson]. Treating the interaction Hamiltonian as a perturbation to the free Hamiltonian is much harder due to unbounded nature of the field operator (see, e.g., [hasler2011ground, hasler2021existence]).
@@
Apart from the issue of mathematical rigor, the main physical reason for choosing Assumption 1 is for the sake of having rigorous results even for the strong coupling regime where the coupling constants $\lambda_{j}$ for each detector can be chosen to be arbitrarily large. In the original UDW case, it was shown that if the interaction Hamiltonian of the UDW model (15) has spacelike supports, then the reduced density matrix of the two detectors after interaction is separable [Simidzija2018no-go]. The key, of course, is spacelike separation as we know that generic interacting quantum systems that are in causal contact will generate some entanglement, and the same gapless model was shown to generate entanglement when they are in causal contact [perche2024closedform].
@@
As we will show in Sec. III, one of our contributions will be to show that not only does the RSB model satisfy the no-go theorem, but that the detectors can only get entangled very deep into the light cone. In other words, even though we can in principle make the final state arbitrarily close to the maximally entangled state, it takes much longer than the light-crossing time between them. To set up the stage for these results, we need more assumptions along the lines used for the single detector case in [Tjoa2024interacting]. These additional assumptions require us to first study the ground state properties of the full detector-field system.
[…]
@@
Proof.
@@
[…]
@@
[…]